OpenStudy (anonymous):

For the function f(x) = –(x + 1)2 + 4, identify the vertex, domain, and range.

OpenStudy (anonymous):

Do you mean \( \huge f(x) = (x-1)^2 + 4) \)

OpenStudy (anonymous):

If you look at \( \huge f(x) = (x-1)^2 + 4) \) you can tell where the vertex is located. \( \huge f(x) = (x-1)^2 + 4) \) The 1 in \( (x-1)^2 \) is the x and the + 4 is the y so the vertex is (1,4)

OpenStudy (anonymous):

It also is telling the parabola to shift to the right by 1 and go up by four so where the 1 is located is known as the horizontal shift area and the 4 is in the vertical shift area.

OpenStudy (anonymous):

What do you think is the domain and range?

OpenStudy (anonymous):

Is the domain all real numbers? And the range is y greater than or equal to 4?

OpenStudy (triciaal):

@Nixy a = -1

OpenStudy (anonymous):

I'm unsure as to whether the vertex is truly (1,4) or if it is (-1,4). For reference, the function is f(x)=-(x+1)^2 +4

OpenStudy (triciaal):